OpenStudy (anonymous):

Find the derivative of f when f(x)= cos(x)-1 / sin(x)

OpenStudy (anonymous):

use quotient rule

OpenStudy (anonymous):

I did, but my answer is still not in the answer choices

OpenStudy (anonymous):

(-sinx)(sinx)- (cosx-1)(cosx) / sin^2x simplify that

OpenStudy (anonymous):

\[(-sinx)(sinx)-(cosx-1)(cosx)\]\[-\sin ^{2}x-\cos ^{2}x+cosx \div \sin ^{2}x\]\[-(\sin ^{2}x + \cos ^{2}x)+ cosx \div \sin ^{2}x\]\[-1+cosx \div \sin ^{2}x\]

OpenStudy (anonymous):

is this correct? because its not in the answer choices...

OpenStudy (anonymous):

\[\frac{ (-sinx)(sinx)-(cosx-1)(cosx) }{ \sin^2x }=\frac{ -(\sin^2x+\cos^2x-cosx) }{ \sin^2x }\]